(5.1) Find the vector form of the equation of the plane that passes through the point P0 = (1, −2, 3) (5)

and has normal vector ~n =< 3, 1, −1 >.

(5.2) Find an equation for the plane that contains the line x = −1 + 3t, y = 5 + 3t, z = 2 + t and is (6)

parallel to the line of intersection of the planes x −2(y −1) + 3z = −1 and y = −2x −1 = 0.


1
Expert's answer
2021-07-07T12:56:35-0400

(5.1) The standard form of the equation of the plane is


"3(x-1)+1(y-(-2))-1(z-3)=0"

"3x+y-z+2=0"

Parameters: "y=s, z=t"



"x=-\dfrac{2}{3}-\dfrac{1}{3}s+\dfrac{1}{3}t"

"y=0+1s+0t"

"z=0+0s+1t"

The vector form of the equation of the plane is


"\vec r=\langle\dfrac{2}{3},0,0\rangle+s\langle-\dfrac{1}{3},1,0\rangle+t\langle\dfrac{1}{3},0,1\rangle"

(5.2)


"x \u22122(y \u22121) + 3z = \u22121""y = \u22122x \u22121"




"y=-2x-1""x+4x+4+3z=-1"

"y=-2x-1""z=-\dfrac{5}{3}x-\dfrac{5}{3}""x\in \R"

The intersection between the planes is the line :


"\langle0,-1,-\dfrac{5}{3}\rangle+t\langle1, -2,-\dfrac{5}{3}\rangle, t\in \R"


The the plane contains the line


"\dfrac{x-0}{1}=\dfrac{y+1}{-2}=\dfrac{z+\dfrac{5}{3}}{-\dfrac{5}{3}}"

An equation for the plane parallel to the line of intersection of the planes


"ax+by+cz+d=0,"

where


"a-2b-\dfrac{5}{3}c=0"

The plane contains the line "x = \u22121 + 3t, y = 5 + 3t, z = 2 + t"

"t=0: Point(-1, 5, 2)"

"t=-2: Point(-7, -1, 0)"


"-a+5b+2c+d=0""-7a-b+d=0"


"b=-7a+d""-a-35a+5d+2c+d=0"


"b=-7a+d""c=18a-3d"

Substitute


"a+14a-2d-30a+5d=0"

"d=5a"

If "a=1"


"d=5"

"b=-2"

"c=3"

The equation for the plane is


"x-2y+3z+5=0,"